how do i solve this question or one similar:
The perimeter of a rectangle is 20cm and its area is 14cm^2. Calculate the dimensions of the rectangle to one decimal place.

Answer by stanbon(75887)

The perimeter of a rectangle is 20cm and its area is 14cm^2. Calculate the dimensions of the rectangle to one decimal place.
------
2(L+W) = 20cm
L*W = 14cm^2
------------------
L+W = 10
----
Substitute for "W" and solve for "L"::
L(10-L) = 14
10L-L^2 = 14
L^2 - 10L + 14 = 0
Use the Quadratic Formula to solve for L::
-------------------
L = 1.68..cm
W = 10-1.68 = 8.32...cm
-------------------------
Cheers,
Stan H.
--------------